a) There are 3 friends Alexis, Kim, and Tim. They always get homework but Alexis does 5 hours less than Kim and Tim does twice the amount of homework as Alexis. Together they have done a total amount 25 hours. Find an equation that represents this statement.
b) Find out how many hours Alexis spends on her homework.
Please help me solve this I'm having a hard time find this out.
Thank you!
2) You earn 75 per/week at your new job. With every thing you sell your boss gives you 4% of the earnings. How much will you get if you sell $1000 in chairs?
3) Your total earning of one week is $125. How many chairs have you sold to earn 125$.
I have this I'm not sure if I'm on the right track.
E=75n+c E=75n+0.04c
125= 525+c 125= 525+0.04c
-525=-525. -525=-525
-400=c -400=0.04c
Check. -10,000=c
125= 525-400 Check
525+0.04 (-10,000) =125
please help

Answer by stanbon(75887)

a) There are 3 friends Alexis, Kim, and Tim. They always get homework but Alexis does 5 hours less than Kim and Tim does twice the amount of homework as Alexis. Together they have done a total amount 25 hours. Find an equation that represents this statement.
-------
a = k-5
t = 2a
a + k + t = 25 hrs
---------------------------
Substitute and solve for "a":
a + a+5 + 2a = 25
4a = 20
a = 5 hrs
b) Find out how many hours Alexis spends on her homework.
Ans: 5 hrs
-----------------------
Please help me solve this I'm having a hard time find this out.
Thank you!
2) You earn 75 per/week at your new job. With every thing you sell your boss gives you 4% of the earnings. How much will you get if you sell $1000 in chairs?Ans: You earn 75 + 0.04*1000 = 75 + 40 = $115.00
----------------------------------------------
3) Your total earning of one week is $125. How many chairs have you sold to earn 125$.
Equation::
125 = 75 + 0.04*sales
50 = 0.04s
sales = 5000/4 = $1250
--------------------------
Cheers,
Stan H.
-----------------